a car at rest is acted upon by a constant force acquiring a velocity of 22 m/s after 4 s . determine the magnitude of constant force if the car weights 10,800 N
1
Expert's answer
2013-09-23T10:52:40-0400
the acceleration is

a=v/t = 22 m/s / 4s = 5.4 m/s^2

The magnitude of constant force is

F = a*m = 5.4*1080 kg = 5832 N.